# A Statistician’s Overview of Mechanistic-Informed Modeling

### v1

@@ 3.1.1 Operational Variational Data Assimilation (4DVar)
In practice, variational data assimilation evolved away from Lagrange multipliers to an unconstrained approach known as “4DVar” (“4D” because it includes 3 dimensions in space and time). As with the Sasaki approach, we have data in the interval $[1,T]$ and seek the initial condition, $\tilde{\bm{u}}_{0}$. The observation component of the objective function is the same as (3) above where again the $\tilde{\bm{u}}_{t}$ are functions of the initial condition $\tilde{\bm{u}}_{0}$ and obtained deterministically from the numerical solution of the PDE (2), which is assumed to be error free. In this approach, one then assumes that there is some background “estimate” of the initial condition, say $\bm{u}_{b}$, such as a forecast from the previous assimilation cycle, or some historical average. Then, the background term of the objective function is given by,
$J_{1b}(\tilde{\bm{u}}_{0})=\frac{1}{2}(\tilde{\bm{u}}_{0}-\bm{u}_{b})^{\prime}\bm{C}_{b_{0}}^{-1}(\tilde{\bm{u}}_{0}-\bm{u}_{b}),$ (5)
where $\bm{C}_{b_{0}}$ is the background error covariance matrix (a spatial covariance matrix for the errors between the true initial state and the previous forecast of the initial state). The optimization minimizes the sum of $J_{0}+J_{1b}$ (i.e., equation (3) plus (5)) . Importantly, including the dynamical constraints implicitly through the deterministic model integration effectively converts the constrained minimization problem (as in (4)) to an unconstrained minimization problem. In practice, the optimization is facilitated by using an adjoint method (e.g., Lewis & Derber,, 1985; Errico,, 1997).
@@ 3.1.2 Weak Constraint 4DVar
One can consider weak-constraint variational assimilation by incorporating an explicit model error term in the objective function (e.g. Bennett,, 1992, via the method of representors). For example, say we have the model error term, $\bm{\eta}_{t}$, as in
$\bm{u}_{t+1}=\mathcal{M}(\bm{u}_{t};\bm{\theta}_{m})+\bm{\eta}_{t},$ (6)
where $\bm{\eta}_{t}\sim\text{Gau}(\bm{\mu}_{\eta},\bm{Q})$, where the mean $\bm{\mu}_{\eta}$ vector and error covariance matrix $\bm{Q}$ are assume known or estimated offline. In this section, we prefer $\bm{u}_{t}$ instead of $\bm{\tilde{u}}_{t}$ since $\bm{u}_{t}$ is not constrained to follow the numerical model exactly, as in the previous subsections. Then, the extra term in the objective function in the 4DVar case is given by (e.g. Laloyaux et al.,, 2020),
$J_{2}(\bm{u}_{0})=\frac{1}{2}\sum_{t=1}^{T}(\bm{\eta}_{t}-\bm{\mu}_{\eta})^{\prime}\bm{Q}^{-1}(\bm{\eta}_{t}-\bm{\mu}_{\eta}),$ (7)
where $\bm{\eta}_{t}\equiv\bm{u}_{t+1}-\mathcal{M}(\bm{u}_{t};\bm{\theta}_{m})$ from (6) and again, $\bm{u}_{0}$ is implicitly included in each of the model error terms, $\bm{\eta}_{t}$. Minimizing the loss function $J_{0}+J_{1b}+J_{2}$ (i.e., equation (3) plus (5) plus (7)) in terms of $\bm{u}_{0}$ is constrained by the numerical model but can deviate from an exact solution of the model, resulting in a weak constraint.
@@ 3.1.3 Optimal Interpolation
Around the same time that the variational approach to meteorological data assimilation was being developed, other researchers were developing statistical optimization approaches based on stochastic processes. A brief history of these approaches is presented in Lewis & Lakshmivarahan, (2008). Although there is some debate about its origins, Lewis & Lakshmivarahan, (2008) describe the first practical approach to come from A. Eliassen in an obscure technical report from Oslo University in 1954. More formal development was given by Gandin, (1963). Gandin’s approach is widely considered to have been equivalent to the formalization of kriging by Matheron (Matheron,, 1963), which is the essence of geostatistical modeling (“kriging”) in spatial statistics; see Cressie, (1990) for a discussion of the origins of kriging. In retrospect, all these approaches simply use Gaussian processes to perform interpolation (spatial prediction) that is optimal in the sense of minimizing the expected squared error loss (i.e., risk).
In the context of meteorological data assimilation, one interpolates the residuals by removing a background field (e.g., the previous forecast) from observations. So, given an observation vector, $\bm{z}_{t}$, and a forecast from the previous time, $\bm{u}_{b}$ (where again, $b$ stands for “background”; note, keeping with the notation used in this paper, when this forecast is from the numerical solution of a PDE, we could denote the background as $\tilde{\bm{u}}_{b}$), the Eliassen (simple kriging) approach gives the latent state vector ${\bm{u}}_{t}$ from:
@@
where $\bm{C}_{b}$ is the $n\times n$ spatial covariance matrix of the background (forecast) error. This matrix must be specified in terms of a covariance function, and because one may need to develop predictions from locations that are not observed, it comes from a covariance function associated with a Gaussian process (e.g., see Cressie & Wikle,, 2011).
There are different ways that physical constraints are included in this framework. First, note that the forecast for time $t$, which we are denoting by $\bm{u}_{b}$ here, comes from a forecast given by the dynamic model (2) and the previous estimate of the state, say ${\bm{u}}_{t-1}$. Alternatively, one can specify covariance functions that satisfy dynamical constraints, which is especially useful in multivariate applications of this methodology (e.g., see Daley,, 1993).
Note, this formulation does not explicity account for multiple observations over times; i.e., there is no dependence specified across time. In addition, this optimal interpolation approach for spatial fields can also be formulated from a variational perspective as in Section 3.1.1, known as “3DVar” in the data assimilation literature. Spatial optimal interpolation and 3DVar are equivalent if the observation mapping function is linear, i.e., $\bm{h}(\bm{u}_{t})=\bm{H}_{t}\bm{u}_{t}$.
@@ 3.1.4 Connections to Filtering and Bayesian Estimation
About the same time that Sasaki was developing the variational approach to combining data and mechanistic model information, control engineers were working on the same problem, from an online filtering perspective. The classic example of this is the Kalman filter (Kalman,, 1960), which has become ubiquitous in control engineering and statistical state-space time-series and spatio-temporal analysis (e.g., see Shumway & Stoffer,, 2000; Wikle & Berliner,, 2007). Typically in the control engineering case, one assumes they have an initial condition and initial variance/covariance matrix and proceeds to generate forecasts and update those as new observations come online, then generates another forecast, etc. The filtering equations explicitly account for model error, as in the weak constraint 4DVar case. Indeed, there is a clear connection between the filtering formula and the variational equations (e.g., see Wikle & Berliner,, 2007, for a derivation and connections).
As discussed in Wikle & Berliner, (2007), one can view the variational, optimal interpolation, and filtering approaches as Bayesian inference, where there is a likelihood model relating observations to a latent dynamic process, and then the process is assigned a prior distribution with the background vector (e.g., forecast) $\bm{u}_{b}$ as the prior mean. When the data and prior distributions are Gaussian and the mapping function is linear, the variational and Bayesian approaches are equivalent (e.g., Wikle & Berliner,, 2007).
@@
One of the limitations of this perspective is that it assumes the model parameters (and error variances/covariances) are known (or they are somehow estimated off-line). In the case of online filters, one can augment the state vector with parameters and update them along with the state process. Alternatively, in variational approaches it is certainly possible to add another term to the objective function that accounts for parameter uncertainty. This is equivalent to assigning a prior distribution to the parameters. When practical, it is often more flexible to specify prior distributions explicitly in case non-Gaussian priors are warranted. The development of mechanistically-motivated spatio-temporal dynamic models (described in the next section) evolved from this hierarchical Bayesian perspective.
@@ 3.2 Statistical Approaches for Adding Mechanistic Information
There have been several approaches to combine mechanistic information in statistical spatial and spatio-temporal models. For example, the development of spatial and spatio-temporal covariance functions from analytical or numerical solutions to PDEs has long been considered an efficient way of incorporating mechanistic information in spatial and spatio-temporal models. PDE-derived covariance functions have been used for spatial prediction, such as for kriging models (e.g., Heine,, 1955; Whittle,, 1954), and have recently become quite popular with the stochastic-PDE (SPDE) approach (e.g., Lindgren et al.,, 2011, 2022). Indeed, as mentioned in Section 3.1.3, this idea of using PDE-derived covariance functions for optimal interpolation has also been used in the early data assimilation literature (e.g., Daley,, 1993). Generally, this approach is useful for developing covariance functions, but is quite limited in the complexity of the underlying mechanistic model that can be effectively considered (see the overview in Cressie & Wikle,, 2011).
As an alternative, in the functional data and non-parametric regression analysis literature, it has long been common to use function derivatives to regularize a regression model such as in smoothing splines for spatial data (e.g., Wahba,, 1990; Eubank,, 1999; Ramsay,, 2002). Over the last 10 years or so, this has been extended to the physics-informed regularization case in spatial and spatio-temporal regression (e.g., see the review in Sangalli,, 2021). These methods are very much in the spirit of the Sasaki, (1958) approach described in Section 3.1, where one adds a PDE regularization term to the loss function that compares the non-parametric model of interest to the data. As described in Tomasetto et al., (2024), point estimates of parameters can be estimated in this framework. These approaches can be quite effective for certain problems and in fundamental ways, the PINN approach described below in Section 4.1 can be thought of as a generalization of this approach with more flexible semi-parametric (deep neural) models and the use of automatic differentiation to allow more complex PDE regularization. Bayesian-PINNs (described in Section 4.2), mechanistically-motivated statistical models (described in Section 3.2.1), and probabilistic numerical methods (described in Section 3.2.2) operate similarly, but can accommodate more principled uncertainty quantification across all phases of model development via a hierarchical Bayesian framework. Given uncertainty quantification is a major focus for modern spatio-temporal modeling applications, we describe the statistical approaches in more detail below in the following subsections, and PINNs and B-PINNs in Section 4.
@@ 3.2.1 Mechanistically-Motivated (Physical-Statistical) Models
Before introducing the mechanistically-motivated framework, we first briefly review Bayesian hierarchical models for spatio-temporal processes. Adding additional sources of data and accounting for uncertainty in observation model parameters and state-process model parameters is often needed when modeling complex spatio-temporal data. In some cases, we can gain much more expressiveness in our models if we add dependence structure (time, space, covariates) to the parameters as well (i.e., treat them as dependent random processes, conditionally). Bayesian hierarchical models are natural extensions of the Bayesian paradigm that allow for more complete accommodation of uncertainty and additional model flexibility.
The Bayesian hierarchical (deep) modeling paradigm for spatio-temporal data (e.g., Berliner,, 1996; Cressie & Wikle,, 2011) considers breaking the joint distribution of data, process, and parameters into a series of conditional models (distributions). In that context, the posterior distribution is then given by the following decomposition
@@
×$\times\;\;\;[process|process\;parameters]$
×$\times\;\;\;[parameters].$
@@
In the context of the dynamic spatio-temporal models (DSTMs) of interest here, a basic BHM might take the following form.
Data Model: $\;\;\;\bm{z}_{t}=\bm{h}_{t}(\bm{u}_{t})+\bm{\epsilon}_{t},\;\;\bm{\epsilon}_{t}\overset{ind.}{\sim}\;\text{Gau}(\bm{0},\bm{R}_{t}(\bm{\theta}_{z})),$
Process (Dynamic) Model: $\;\;\;\bm{u}_{t}=\mathcal{M}(\bm{u}_{t-1},\bm{B}(\bm{u}_{t};\bm{\theta}_{b});\bm{\theta}_{m})+\bm{\eta}_{t},\;\;\bm{\eta}_{t}\overset{ind.}{\sim}\;\text{Gau}(\bm{0},\bm{Q}_{t}(\bm{\theta}_{q})),$
Process Initial Value: $\;\;\;\bm{u}_{0}\sim\text{Gau}(\bm{a},\bm{\Sigma}_{0})$
@@
where $\bm{B}(\bm{u}_{t};\bm{\theta}_{b})$ represents a boundary condition and $\bm{u}_{0}$ is the initial condition, both of which depend on parameters that are given prior distributions in general. Importantly, the parameter models can also have dependence (in space and/or time) or depend on other covariates, which gives this model a great amount of generalizability and can make it much “deeper” in terms of having more levels of conditioning.
@@
Data: A strength of the BHM approach is that one can easily specify multiple data sources conditioned on the latent state process, e.g., $[Z_{1},Z_{2}|u]=[Z_{1}|u][Z_{2}|u]$ (assuming conditional independence). In statistics, using mechanistic model output (e.g., weather reanalysis fields) as “data” (say, $Z_{2}$) has been a common “data fusion” approach to add mechanistic information since the late 1990s (e.g., Berliner et al.,, 1999; Wikle et al.,, 2001; Berliner,, 2003; Fuentes & Raftery,, 2005; Wikle & Berliner,, 2005). This is a type of weak constraint mechanistic modeling that has the advantage of being able to account for additive errors (to facilitate uncertainty quantification).
@@
Process: As with the data assimilation literature, one can specify $\mathcal{M}(\cdot)$ to be a known mechanistic model. Alternatively, in the physical-statistical (mechanistically-motivated) DSTM paradigm, one can specify statistical models, $\bm{M}(\cdot;\bm{\theta}_{m})$, that are “motivated” by mechanistic processes, but are conditioned on parameters ($\bm{\theta}_{m}$) that typically have spatial or temporal structure (e.g. spatially-varying diffusion parameters); see Wikle & Hooten, (2010) and Cressie & Wikle, (2011) for numerous examples. For example, one may discretize (or write in spectral/finite-element space) a reaction-diffusion PDE or integro-difference equation and condition this on the diffusion/growth parameters that may vary with space as a GP, or in terms of covariates, which are specified at the next level of the hierarchy (e.g., Wikle,, 2003; Xu et al.,, 2005; Hooten & Wikle,, 2008). If desired, one can include hard constraints (i.e., “mass balance”) in the physical-statistical framework (e.g. Berliner et al.,, 2008; Zhuang & Cressie,, 2014; Gopalan et al.,, 2018).
@@
Parameters: In some cases, mechanistic information can be imparted to the parameters (say, $\bm{\theta}_{m}$) through the prior specification; e.g., regularization or informed by mechanistic information. For example, Wikle, (2003) allowed covariates and a latent Gaussian process to model the spatially-varying diffusion rate in reaction-diffusion model for spread of an invasive species and Wikle et al., (2001) included a known turbulent scaling relationship as a prior on the variances of tropical surface winds to add realistic structure. Finally, Leeds et al., (2014) used pre-training/meta-learning of prior distributions with model output obtained from a coupled ocean-ecosystem model to inform a quadratic nonlinear mechanistically-motivated statistical model for data assimilation of ocean color measurements.

### v2

@@ 3.1.1 Operational Variational Data Assimilation (4DVar)
In practice, variational data assimilation evolved away from Lagrange multipliers to an unconstrained approach known as “4DVar” (“4D” because it includes 3 dimensions in space and time). As with the Sasaki approach, one considers data in the interval $[1,T]$ and seeks the initial condition, $\bm{v}_{0}$. The observation component of the objective function is the same as (3) above where again the $\bm{v}_{t}$ are functions of the initial condition $\bm{v}_{0}$ and obtained deterministically from the numerical solution of the PDE (2), which is assumed to be error free. In this approach, one then assumes that there is some background “estimate” of the initial state condition, say $\bm{v}_{b}$, such as a forecast from the previous assimilation cycle, or some historical average. Then, the background term of the objective function is given by,
$J_{1b}(\bm{v}_{0})=\frac{1}{2}(\bm{v}_{0}-\bm{v}_{b})^{\prime}\bm{C}_{b_{0}}^{-1}(\bm{v}_{0}-\bm{v}_{b}),$ (4)
where $\bm{C}_{b_{0}}$ is the background error covariance matrix (a spatial covariance matrix for the errors between the true initial state and the previous forecast of the initial state). The optimization finds $\bm{v}_{0}$ that minimizes the sum of $J_{0}(\bm{v}_{0})+J_{1b}(\bm{v}_{0})$ (i.e., equation (3) plus (4)). In practice, the optimization is facilitated by using an adjoint method (e.g., lewis1985use; errico1997adjoint).
@@ 3.1.2 Weak Constraint 4DVar
One can consider weak-constraint variational assimilation by incorporating an explicit model error term in the objective function (e.g., bennett1992inverse, via the method of representors). For example, say we have the model error term, $\bm{\eta}_{t}$, as in
$\bm{u}_{t}=\mathcal{M}(\bm{u}_{t-1};\bm{\theta}_{m})+\bm{\eta}_{t},$ (5)
where $\bm{\eta}_{t}\sim\text{N}(\bm{\mu}_{\eta},\bm{Q})$, and the mean $\bm{\mu}_{\eta}$ vector and error covariance matrix $\bm{Q}$ are assumed known or estimated offline. In this section, we prefer to denote the process by $\bm{u}_{t}$ instead of ${\bm{v}}_{t}$ since $\bm{u}_{t}$ is not constrained to follow the numerical model exactly, as in the previous subsections.
Then, the extra term in the objective function in the 4DVar case is given by (e.g. laloyaux2020exploring),
$J_{2}(\bm{\eta})=\frac{1}{2}\sum_{t=1}^{T}(\bm{\eta}_{t}-\bm{\mu}_{\eta})^{\prime}\bm{Q}^{-1}(\bm{\eta}_{t}-\bm{\mu}_{\eta}),$ (6)
where $\bm{\eta}_{t}\equiv\bm{u}_{t}-\mathcal{M}(\bm{u}_{t-1};\bm{\theta}_{m})$ from (5) and $\bm{\eta}=\{\bm{\eta_{1}},...,\bm{\eta_{T}}\}$. Minimizing the loss function $J_{0}(\bm{u_{0}},\bm{\eta})+J_{1b}(\bm{u_{0})}+J_{2}(\bm{\eta})$ (i.e., equation (3) plus (4) plus (6)) in terms of $\bm{u}_{0}$ and $\bm{\eta}$ is constrained by the numerical model but can deviate from an exact solution of the model because of $\bm{\eta}$, resulting in a weak constraint. Note that $J_{0}$ is written as a function of both $\bm{u}_{0}$ and $\bm{\eta}$ because both are necessary to determine $\bm{u}_{1},...,\bm{u}_{T}$ as per (5).
@@ 3.1.3 Optimal Interpolation
Around the same time that the variational approach to meteorological data assimilation was being developed, other researchers were developing statistical optimization approaches based on stochastic processes. A brief history of these approaches is presented in lewis2008sasaki. Although there is some debate about its origins, lewis2008sasaki describe the first practical approach to come from eliassen1954provisional. More formal development was given by gandin1963objective. Gandin’s approach is widely considered to have been equivalent to the formalization of kriging by Matheron (matheron1963principles), which is the essence of geostatistical modeling (“kriging”) in spatial statistics; see cressie1990origins for a discussion of the origins of kriging. In retrospect, all these approaches simply use Gaussian processes to perform interpolation (spatial prediction) that is optimal in the sense of minimizing the expected squared error loss (i.e., risk).
In the context of meteorological data assimilation, one interpolates the residuals by removing a background field (e.g., the previous forecast) from observations. So, given an observation vector, $\bm{z}_{t}$, and a forecast from the previous time, $\bm{u}_{b}$ (where again, $b$ stands for “background”; note, keeping with the notation used in this paper, when this forecast is from the numerical solution of a PDE, we could denote the background as ${\bm{v}}_{b}$), the kriging approach gives the latent state vector ${\bm{u}}_{t}$ from:
@@
where $\bm{C}_{b}$ is the $n\times n$ spatial covariance matrix of the background (forecast) error. This matrix must be specified in terms of a covariance function, and because one may need to develop predictions from locations that are not observed, it comes from a covariance function associated with a Gaussian process (e.g., see cressie1993statistics; cressie2011statistics).
There are different ways that physical constraints are included in this framework. First, note that the forecast for time $t$, which we are denoting by $\bm{u}_{b}$ here, comes from a forecast given by the dynamic model (2) and the previous estimate of the state, say ${\bm{u}}_{t-1}$. Alternatively, one can specify covariance functions that satisfy dynamical constraints, which is especially useful in multivariate applications of this methodology (e.g., see daley1993atmospheric).
Note, this formulation does not explicity account for multiple observations over times; i.e., there is no dependence specified across time. In addition, this optimal interpolation approach for spatial fields can also be formulated from a variational perspective as in Section 3.1.1, known as “3DVar” in the data assimilation literature. Spatial optimal interpolation and 3DVar are equivalent if the observation mapping function is linear, i.e., $\bm{h}(\bm{u}_{t})=\bm{H}_{t}\bm{u}_{t}$ (e.g., lewis2008sasaki).
@@ 3.1.4 Connections to Filtering and Bayesian Estimation
About the same time that Sasaki was developing the variational approach to combining data and mechanistic model information, control engineers were working on the same problem from an online filtering perspective. The classic example of this is the Kalman filter (kalman1960new), which has become ubiquitous in control engineering and statistical state-space time-series and spatio-temporal analysis (e.g., see shumway2000time; wikle2007bayesian). In the control engineering case, one typically assumes they have an initial condition and initial variance/covariance matrix and proceeds to generate a forecast for the next time; this forecast is then updated as new observations come online, and this updated state estimate is used to generate another forecast, etc. The filtering equations explicitly account for model error, as in the weak constraint 4DVar case.
It has long been known that there are connections between the variational optimization approach in data assimilation and Bayesian inference (e.g., lorenc1986analysis; talagrand1997assimilation; wikle2007bayesian). This is not surprising to statisticians, as the maximum a posteriori (MAP) estimator in a Bayesian model is equivalent to a regularized optimization. That is, say we have data $\bm{D}$ and parameters $\bm{\theta}$, then the MAP estimator is given by maximizing the numerator in Bayes formula with respect to $\bm{\theta}$
$\hat{\bm{\theta}}_{MAP}$ $=$ $\underset{\bm{\theta}}{\arg\max}([\bm{D}|\bm{\theta}][\bm{\theta}])$
$=$ $\underset{\bm{\theta}}{\arg\max}(\log[\bm{D}|\bm{\theta}]+\log[\bm{\theta}]),$
which of course is equivalent to minimizing $(-\log[\bm{D}|\bm{\theta}]-\log[\bm{\theta}])$. The connection is then obvious in that the first term (Bayesian likelihood) corresponds to a loss function such as (3) and the second term (Bayesian prior) corresponds to a regularization term such as (4). A classic example in statistics is ridge regression, where a least squares regression is regularized by an $\ell_{2}$ penalty on the parameters, which is equivalent to a multivariate normal prior on the parameters. Typically, assuming one has a reasonably strong belief that their models are correct, the Bayesian approach is often preferred because one gets natural uncertainty quantification. However, in the context of high-dimensional nonlinear data and process models that are used in operational data assimilation, the variational approach has been an effective and reasonably efficient assimilation approach where fully Bayesian estimation would be computationally prohibitive.
Similarly, the sequential estimation in the Kalman filter procedure is also naturally considered from a sequential Bayesian perspective (e.g., talagrand1997assimilation; wikle2007bayesian). That is, the forecast distribution for time $t+1$ can be considered as the “prior” distribution for updating the latent state for time $t+1$ when new data become available for that time. For linear, normal data and process models, these update equations are available in closed form and are equivalent to the variational approach, but when these distributions are non-normal or the dynamics are nonlinear, then sequential Monte Carlo approaches such as particle filtering or ensemble Kalman filtering must be used (e.g., see the tutorial in katzfuss2016understanding).
@@
One of the limitations of considering the equivalence of the variational and Bayesian perspective is that it assumes the model parameters (and error variances/covariances) are known (or they are somehow estimated off-line) and that one’s models are correct. In the case of online sequential filtering, one can augment the state vector with parameters and update them along with the state process. Alternatively, in variational approaches it is certainly possible to add another term to the objective function that accounts for parameter uncertainty. This is equivalent to assigning a prior distribution to the parameters. When practical, it is often more flexible to specify prior distributions explicitly in case non-normal priors are warranted. The development of mechanistically-motivated spatio-temporal dynamic models (described in the next section) evolved from this hierarchical Bayesian perspective.
@@ 3.2 Statistical Approaches for Adding Mechanistic Information
There have been several approaches to combine mechanistic information in statistical spatial and spatio-temporal models. For example, the development of spatial and spatio-temporal covariance functions from analytical or numerical solutions to PDEs has long been considered an efficient way of incorporating mechanistic information in spatial and spatio-temporal models. PDE-derived covariance functions have been used for spatial prediction, such as for kriging models (e.g., heine1955models; whittle1954stationary), and have recently become quite popular with the stochastic-PDE (SPDE) approach (e.g., lindgren2011explicit; lindgren2022spde). Indeed, as mentioned in Section 3.1.3, this idea of using PDE-derived covariance functions for optimal interpolation has also been used in the early data assimilation literature (e.g., daley1993atmospheric). Generally, this approach is useful for developing covariance functions, but is quite limited in the complexity of the underlying mechanistic model that can be effectively considered (see the overview in cressie2011statistics).
As an alternative, in the functional data and non-parametric regression analysis literature, it has long been common to use function derivatives to regularize a regression model such as in smoothing splines for spatial data (e.g., wahba1990spline; eubank1999nonparametric; ramsay2002spline). Over the last 10 years or so, this has been extended to the physics-informed regularization case in spatial and spatio-temporal regression (e.g., see the review in sangalli2021spatial). These methods are very much in the spirit of the sasaki1958objective approach described in Section 3.1, where one adds a PDE regularization term to the loss function that compares the non-parametric model of interest to the data. As described in tomasetto2024modeling, point estimates of parameters can be estimated in this framework. These approaches can be quite effective for certain problems, and in fundamental ways the Physics-Informed Neural Network (PINN) approach described below in Section 4.1 can be thought of as a generalization of this approach with more flexible semi-parametric (deep neural) models and the use of automatic differentiation to allow more complex PDE regularization. Bayesian-PINNs (described in Section 4.2), mechanistically-motivated statistical models (described in Section 3.2.1), and probabilistic numerical methods (described in Section 3.2.2) operate similarly, but can accommodate more principled uncertainty quantification across all phases of model development via a hierarchical Bayesian framework. Given that uncertainty quantification is a major focus for modern spatio-temporal modeling applications, we describe the statistical approaches in more detail in the following subsections, and PINNs and B-PINNs in Section 4.
@@ 3.2.1 Mechanistically-Motivated (Physical-Statistical) Models
Before introducing the mechanistically-motivated framework, we first briefly review Bayesian hierarchical models for spatio-temporal processes. Adding additional sources of data and accounting for uncertainty in observation model parameters and state-process model parameters is often needed when modeling complex spatio-temporal data. In some cases, one can gain much more expressiveness in their models if they add dependence structure (time, space) to the parameters as well by considering them to be spatial or temporal random processes and/or to be related to time or space-varying covariates. Bayesian hierarchical models are natural extensions of the Bayesian paradigm that allow for more complete accommodation of uncertainty and this additional model flexibility.
The Bayesian hierarchical (deep) modeling paradigm for spatio-temporal data (e.g., berliner1996hierarchical; cressie2011statistics) considers breaking the joint distribution of data, process, and parameters into a series of conditional models (distributions). In that context, the posterior distribution is then given by the following general decomposition
@@
$\times\;\;\;[process\;|\;process\;parameters]$
$\times\;\;\;[parameters].$
@@
In the context of the dynamic spatio-temporal models (DSTMs) of interest here, the basic BHM might take the following form (e.g., see cressie2011statistics).
Data Model: $\;\;\;\bm{z}_{t}=\bm{h}_{t}(\bm{u}_{t})+\bm{\epsilon}_{t},\;\;\bm{\epsilon}_{t}\overset{ind.}{\sim}\;\text{N}(\bm{0},\bm{R}_{t}(\bm{\theta}_{z})),$
Process (Dynamic) Model: $\;\;\;\bm{u}_{t}=\mathcal{M}(\bm{u}_{t-1},\bm{B}(\bm{u}_{t};\bm{\theta}_{b});\bm{\theta}_{m})+\bm{\eta}_{t},\;\;\bm{\eta}_{t}\overset{ind.}{\sim}\;\text{N}(\bm{0},\bm{Q}_{t}(\bm{\theta}_{q})),$
Process Initial Value: $\;\;\;\bm{u}_{0}\sim\text{N}(\bm{a},\bm{\Sigma}_{0})$
@@
where $\bm{h}_{t}()$ is a mapping function between the latent process and observations, $\text{N}(\bm{\mu},\bm{\Sigma})$ is a multivariate normal distribution with mean $\bm{\mu}$} and covariance matrix $\bm{\Sigma}$}, and $\bm{B}(\bm{u}_{t};\bm{\theta}_{b})$ represents a boundary condition and $\bm{u}_{0}$ is the initial condition, both of which depend on parameters that are given prior distributions in general. The unknown parameters are given by ${\bm{\theta}}_{z},{\bm{\theta}}_{m},{\bm{a}},{\bm{\Sigma}}_{0},{\bm{\theta}}_{b},{\bm{\theta}}_{q},$. As mentioned above, the parameter models can also have dependence (in space and/or time) or depend on other covariates, which gives this model a great amount of generalizability and can make it much “deeper” in terms of having more levels of conditioning (e.g., see Process description below).
@@
Data: A strength of the BHM approach is that one can easily specify multiple data sources conditioned on the latent state process, e.g., $Z_{1},Z_{2}|u]=[Z_{1}|u][Z_{2}|u]$ (assuming conditional, although this assumption can be relaxed). In statistics, a common “data fusion” approach since the late 1990s has been to combine in situ data, such as satellite observations ($Z_{1}$) and mechanistic model outputs such as weather reanalysis fields ($Z_{2}$), to incorporate mechanistic information (e.g., berliner1999bayesian; wikle2001spatiotemporal; berliner2003physical; fuentes2005model; wikle2005combining). This is a type of weak constraint mechanistic modeling that has the advantage of being able to account for additive errors (to facilitate uncertainty quantification).
@@
Process: As with the data assimilation literature, one can specify $\mathcal{M}(\cdot)$ to be a known mechanistic model. Alternatively, in the physical-statistical (mechanistically-motivated) DSTM paradigm, one can specify statistical models for the transition operator, say $\bm{M}(\cdot;\bm{\theta}_{m})$, that are “motivated” by mechanistic processes, but are conditioned on parameters $\bm{\theta}_{m}$ that typically have spatial or temporal structure (e.g. spatially-varying diffusion parameters); see wikle2010 and cressie2011statistics for numerous examples. For example, one may discretize (or write in spectral/finite-element space) a reaction-diffusion PDE or integro-difference equation and condition this on the diffusion/growth parameters that may vary with space as a GP, or in terms of covariates, which are specified at the next level of the hierarchy (e.g., wikle2003hierarchical; xu2005kernel; wikle2007bayesian; hooten2008).
@@
Parameters: In some cases, mechanistic information can be imparted to the parameters (say, $\bm{\theta}_{m}$) through the prior specification; e.g., via regularization or informed by mechanistic information. For example, wikle2003hierarchical allowed covariates and a latent Gaussian process to model the spatially-varying diffusion rate in reaction-diffusion model for spread of an invasive species and wikle2001spatiotemporal included a known turbulent scaling relationship as a prior on the variances of tropical surface winds to add realistic structure. Alternatively, leeds2014emulator used pre-training/meta-learning of prior distributions with model output obtained from a coupled ocean-ecosystem model to inform a quadratic nonlinear mechanistically-motivated statistical model for data assimilation of ocean color measurements.
As a concrete example illustrating the process and parameter stages of a mechanistically motivated DSTM, consider the latent dynamic process $v$ in one-dimensional space, which is assumed to approximately follow the advection–diffusion equation
$\frac{\partial v}{\partial t}=a\frac{\partial v}{\partial s}+\lambda\frac{\partial^{2}v}{\partial s^{2}}.$ (7)
As shown in wikle2007bayesian, a simple finite difference numerical solution suggests the following difference equation
$v_{t}(s)=\theta_{1}v_{t-\Delta_{t}}(s)+\theta_{2}v_{t-\Delta_{t}}(s+\Delta_{s})+\theta_{3}v_{t-\Delta_{t}}(s-\Delta_{s}),$
where the coefficients $\theta_{i}$ are functions of the spatial discretization, $\Delta_{s}$, time discretization, $\Delta_{t}$, and advection and diffusion parameters, $a$ and $\lambda$, respectively. Denoting the discretized process at $n$ spatial locations by the vector ${\bm{v}}_{t}=[v_{t}(s_{1}),\ldots,v_{t}(s_{n})]^{\prime}$, and the boundary process by ${\bm{v}}_{t}^{b}=[v_{t}(s_{0}),v_{t}(s_{n+1})]^{\prime}$, we can write the vectorized difference equation for the solution of the PDE as ${\bm{v}}_{t}={\bm{M}}{\bm{v}}_{t-\Delta_{t}}+{\bm{M}}^{b}{\bm{v}}^{b}_{t-\Delta_{t}}$, where $\bm{M}$} and ${\bm{M}}^{b}$ are functions of $\theta_{0},\theta_{1},\theta_{2}$; e.g., $\bm{M}$} is a tridiagonal matrix (e.g., with $\theta_{0}$ on the main diagonal and $\theta_{1},\theta_{3}$ on the off-diagonals; e.g., see wikle2007bayesian). Now, a mechanistically motivated model for the latent process of interest, say ${\bm{u}}_{t}$, could be written as ${\bm{u}}_{t}={\bm{M}}_{u}{\bm{u}}_{t-1}+{\bm{\eta}}_{t}$, where ${\bm{M}}_{u}$ might be a tridiagonal matrix, but where the main diagonal is spatially varying, e.g., ${\bm{\theta}}_{0}\sim\text{N}({\bm{X}}{\bm{\beta}},{\bm{\Sigma}}_{\theta})$, where ${\bm{\theta}}_{0}$ is an $n\times 1$ vector and $\bm{X}$ represents covariates and $\bm{\Sigma}_{\theta}$ imparts spatial dependence. With a data model that relates observations to the latent process, ${\bm{u}}_{t}$, this corresponds to a simple version of the BHM DSTM described above. The aformentioned references contain specific examples and extensions of such a mechanistically motivated model. For example, if desired, one can include hard constraints (i.e., “mass balance”) in the physical-statistical framework (e.g. berliner2008modeling; zhuang2014bayesian; gopalan2018bayesian).